# Arbitrarily-Conditioned Multi-Functional Diffusion for Multi-Physics Emulation

### v1

@@ Abstract
Modern physics simulation often involves multiple functions of interests, and traditional numerical approaches are known to be complex and computationally costly. While machine learning-based surrogate models can offer significant cost reductions, most focus on a single task, such as forward prediction, and typically lack uncertainty quantification — an essential component in many applications. To overcome these limitations, we propose Arbitrarily-Conditioned Multi-Functional Diffusion (ACM-FD), a versatile probabilistic surrogate model for multi-physics emulation. ACM-FD can perform a wide range of tasks within a single framework, including forward prediction, various inverse problems, and simulating data for entire systems or subsets of quantities conditioned on others. Specifically, we extend the standard Denoising Diffusion Probabilistic Model (DDPM) for multi-functional generation by modeling noise as Gaussian processes (GP). We then introduce an innovative denoising loss. The training involves randomly sampling the conditioned part and fitting the corresponding predicted noise to zero, enabling ACM-FD to flexibly generate function values conditioned on any other functions or quantities. To enable efficient training and sampling, and to flexibly handle irregularly sampled data, we use GPs to interpolate function samples onto a grid, inducing a Kronecker product structure for efficient computation. We demonstrate the advantages of ACM-FD across several fundamental multi-physics systems.
@@ 1 INTRODUCTION
Physical simulation plays a crucial role in numerous scientific and engineering applications. Traditional numerical approaches (Zienkiewicz et al., 1977; Mitchell and Griffiths, 1980), while offering strong theoretical guarantees, are often complex to implement and computationally expensive to run. In contrast, machine learning-based surrogate models (Kennedy and O’Hagan, 2000; Razavi et al., 2012), also known as emulators, trained on simulation or measurement data, can significantly reduce computational costs, making them a promising alternative.
However, modern physical simulations often involve multiple functions of interest, such as initial and boundary conditions, solution or state functions, parameter functions, source functions, and more. Current machine learning-based surrogate models, such as neural operators (Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2023), primarily focus on a single prediction task, for instance, forward prediction of the solution function. To perform other tasks, one typically needs to retrain a surrogate model from scratch. In addition, most existing models lack uncertainty quantification, which is important in many applications. For example, confidence intervals are important to access the reliability of emulation results, and in inverse problems, a posterior distribution is required since such problems are often ill-posed.
@@
  • Multi-Functional Diffusion Framework: we propose a multi-functional diffusion framework based on the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). By modeling noise as multiple Gaussian processes (GPs), we perform diffusion and denoising in functional spaces, enabling the generation of multiple functions required in multi-physics systems.
  • Innovative Denoising Loss: we introduce a denoising loss that encapsulates all possible conditional parts within the system. During training, we randomly sample the conditioned part and fit the predicted noise for those parts to zero. This enables ACM-FD to flexibly generate function values conditioned on any other set of functions or quantities, allowing it to handle a wide range of tasks, such as forward prediction, inverse inference, and data simulation.
  • Efficient Training and Sampling: To enable efficient training and sampling, and to flexibly handle irregularly sampled data that might come from real-world experiments and measurements, we use GPs to interpolate irregular function samples onto a grid. On the grid, we employ a multiplicative kernel to induce a Kronecker product structure within the kernel matrix. Leveraging the properties of the Kronecker product and tensor algebra, we avoid computing the full kernel matrix and its Cholesky decomposition, thereby substantially reducing the training and sampling costs.
  • Experiments: We evaluated ACM-FD on three fundamental multi-physics systems, including Darcy flow, convection-diffusion, and torus fluid, with the number of involved functions ranging from three to seven. In twenty prediction tasks across the three systems, ACM-FD consistently achieved top-tier performance compared to state-of-the-art neural operators specifically trained for each task. Furthermore, we evaluated ACM-FD in data simulation. We compared with a popular variational auto-encoder approach. The data generated by ACM-FD not only adheres much more closely to the governing equations but also exhibits substantially greater diversity.
@@ 2 PRELIMINARIES
where $\widehat{\beta}_{t}=\frac{1-\widehat{\alpha}_{t-1}}{1-\widehat{\alpha}_{t}}\beta_{t}$. Accordingly, DDPM employs a neural network $\Phi$ to approximate the noises $\bm{\xi}_{t}$; see (2) The training objective is to minimize the denoising loss, $\mathbb{E}_{t}\|\Phi_{\Theta}({\bf x}_{t},t)-\bm{\xi}_{t}\|^{2}$.
@@ 3 METHOD
We consider a multi-physics system involving a collection of $M$ functions. These functions may represent initial and boundary conditions, source functions, parameter functions, system states and others. Our goal is to capture the underlying complex relationships between these functions and perform a wide range of prediction and simulation tasks, such as forward prediction of the system state given the initial condition, inverse inference about the systems parameters given observed states, joint simulation of all the functions, or conditional simulation of one set of functions based on another.
Many of these tasks involve mapping between functions, making neural operators (Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2023) — a recently developed class of surrogate models — an appealing approach. However, most neural operators are trained for a single prediction task. To perform a different task, one would need to train another model from scratch. Additionally, these models generally lack uncertainty quantification, which is important for many applications. Furthermore, these models are unable to perform data generation tasks.
To address these issues, we propose AFML, an innovative generative model that can serve as a powerful and versatile probabilistic emulator for multi-physics systems.
@@ 3.1 Multi-Functional Diffusion
where $f_{t}(\cdot)$ is the noisy version of $f_{0}$ at step $t$, and $\xi_{t}(\cdot)$ is the noise function used to corrupt $f_{0}$. To model the noise function, one can employ an expressive stochastic process. We choose to sample each noise function from a zero-mean Gaussian process (GP) (Rasmussen and Williams, 2006),
@@
where $\kappa(\cdot,\cdot)$ is the covariance (kernel) function and ${\bf z}$ and ${\bf z}^{\prime}$ denote the input locations of the function. When $t$ becomes sufficiently large (i.e., $\widehat{\alpha}_{t}\approx 0$), we can approximate $f_{t}\sim\mathcal{GP}(0,\kappa({\bf z},{\bf z}^{\prime}))$, meaning it effectively turns into a noise function.
The actual data instance is the function sampled at a finite set of locations, ${\bf f}_{0}=f_{0}({\mathcal{Z}})=\left(f_{0}({\bf z}_{1}),\ldots,f_{0}({\bf z}_{N})\right)^{\top}$ where ${\mathcal{Z}}=\{{\bf z}_{j}|1\leq j\leq N\}$ are the sampling locations. Therefore, we only need to apply the diffusion process (5) to ${\mathcal{Z}}$ (with the function values at all the other locations marginalized out), which yields:
@@
Comparing to (4), the only difference is that the identity matrix is replaced by the covariance matrix ${\bf K}$ at the sampling locations ${\mathcal{Z}}$. We then use a similar strategy as in DDPM. We train a neural network $\Phi_{\Theta}$ such that $\Phi_{\Theta}({\bf f}_{t},t,{\mathcal{Z}})\approx\bm{\xi}_{t}$, and substitute the neural network’s prediction for $\bm{\xi}_{t}$ in (8) to obtain $p_{\Theta}({\bf x}_{t-1}|{\bf x}_{t})$ for denoising and data sampling.
We next generalize to the multi-functional case. Given $M$ functions of interest, =$\mathcal{F}=\{f^{1}_{0}(\cdot),\ldots,f^{M}_{0}(\cdot)\}$, we employ the same diffusion process as specified in (5) and (7) to transform each function into a noise function. Next, we use a single neural network to jointly recover or sample of all the functions, thereby capturing the complex and strong relationships among them. Specifically, we construct a network $\Phi_{\Theta}$ that takes the step $t$, all the noisy function values sampled at $t$, and the sampling locations as the input, to predict the corresponding noises,
@@
where each ${\bf f}^{k}_{t}=f^{k}_{t}({\mathcal{Z}}_{k})$, $\bm{\xi}^{k}_{t}=\xi^{k}_{t}({\mathcal{Z}}_{k})$, $f^{k}_{t}(\cdot)$ is the noisy version of $f^{k}_{0}(\cdot)$ at step $t$, $\xi^{k}_{t}(\cdot)$ is the corresponding noise function that corrupts $f^{k}_{0}$ to yield $f^{k}_{t}$, and ${\mathcal{Z}}_{k}$ are the sampling locations of $f^{k}_{0}$, and $\overline{{\mathcal{Z}}}=\{{\mathcal{Z}}_{k}\}_{k=1}^{M}$.
@@ 3.2 Arbitrarily-Conditioned Denoising Loss
Specifically, let us denote an arbitrary set of conditioned function values as ${\bf F}^{c}=\{f_{0}^{k}({\mathcal{Z}}_{k}^{c})|k\in\mathcal{C}\}$, and the target function values to be generated as ${\bf F}^{s}=\{f_{0}^{k}({\mathcal{Z}}_{k}^{s})|k\in{\mathcal{S}}\}$. Note that $\mathcal{C}$ and ${\mathcal{S}}$ may overlap, provided that for any $f^{k}\in\mathcal{C}\cap{\mathcal{S}}$, the sampling locations do not, i.e., ${\mathcal{Z}}_{k}^{c}\cap{\mathcal{Z}}_{k}^{s}=\emptyset$. To generate ${\bf F}^{s}$ conditioned on ${\bf F}^{c}$, we fix ${\bf F}_{c}$ in the input to the network $\Phi_{\Theta}$, and vary only the noisy state associated with ${\bf F}_{s}$, denoted as ${\bf F}^{s}_{t}$. The input to $\Phi$ consists of ${\bf F}^{c}\cup{\bf F}^{s}_{t}\cup\{t,\overline{{\mathcal{Z}}}\}$. We use the predicted noise corresponding to ${\bf F}^{s}_{t}$ to sample ${\bf F}^{s}_{t-1}$, following the denoising process. This procedure is repeated until $t=0$, resulting in a sample for ${\bf F}^{s}$. The sampling process is summarized in Algorithm 2.
@@
By varying the choices of ${\bf F}^{c}$ and ${\bf F}^{s}$, our model can perform a wide variety of data generation and prediction tasks. When ${\bf F}^{c}=\emptyset$, the model jointly samples the $M$ target functions at the specified input locations. When ${\bf F}^{c}\neq\emptyset$ and $\mathcal{C}\cap{\mathcal{S}}=\emptyset$, the model generates one set of functions conditioned on the other, supporting various forward prediction and inverse inference tasks. For instance, by setting $\mathcal{C}$ to include initial conditions or current and/or past system states and source functions, and ${\mathcal{S}}$ to the future system states, the model performs forward prediction. Alternatively, if the observed system states are included in $\mathcal{C}$ and ${\mathcal{S}}$ corresponds to initial/boundary conditions, system parameters, or past states, the model carries out inverse inference. Even for the same type of tasks, the functions and the number of functions in $\mathcal{C}$, as well as the number of sampling locations in ${\mathcal{Z}}_{k}^{c}$, can be adjusted to accommodate different levels of available information for prediction. When $\mathcal{C}\cap{\mathcal{S}}\neq\emptyset$, the model performs not only conditional sampling of other functions but also completion of the observed functions at new locations. All tasks are executed using a single model, $\Phi_{\Theta}$, following the same denoising process. One can generate as many predictive samples as needed, naturally supporting uncertainty quantification in any context. Thereby, our method offers a versatile probabilistic surrogate for multi-physics emulation.
@@
However, to build a well-performing model, training $\Phi$ to predict all the noises as specified in (9) is far from sufficient, because it only handles unconditioned sampling. To make our model capable of handing all kinds of tasks, we introduce a random mask variable, =${\mathcal{H}}=\{{\bf h}^{1},\ldots,{\bf h}^{M}\}$. Each element in ${\mathcal{H}}$ is binary, indicating whether the corresponding function value is conditioned or to be sampled. If a function value is conditioned, then it should remain fixed as a constant input to $\Phi$, and correspondingly, the predicted noise for that value should be zero, since it is never corrupted by any noise. Accordingly, we introduce a new denoising loss,
@@
Here ${\bf f}^{k}_{0}$ is the original function instance sampled at the specified locations ${\mathcal{Z}}^{k}$, ${\bf f}^{k}_{t}=\sqrt{\widehat{\alpha}_{t}}{\bf f}^{k}_{0}+\sqrt{1-\widehat{\alpha}_{t}}\bm{\xi}^{k}_{t}$ is the noisy version of ${\bf f}^{k}_{0}$ at step $t$, and $\circ$ denotes the element-wise product. The choice of $p({\mathcal{H}})$ is flexible. Without any prior knowledge or preferences, we can simply set $p({\mathcal{H}})=\prod_{h\in{\mathcal{H}}}\text{Bernoulli}(h|0.5)$. Our denoising loss (10) accommodates all possible conditional components for conditioned sampling as well as unconditioned sampling (i.e., =${\mathcal{H}}=\{0\}$), allowing our model to be trained for all kinds of tasks. We use stochastic training, which involves randomly sampling the mask ${\mathcal{H}}$, applying it to the function and noise instances as specified in (11), and computing the stochastic gradient $\nabla_{\Theta}\left\|\Phi_{\Theta}(\widehat{{{\bf f}}}^{1}_{t},\ldots,\widehat{{{\bf f}}}^{M}_{t},t,\overline{{{\mathcal{Z}}}})-(\widehat{{\bm{\xi}}}^{1}_{t},\ldots,\widehat{{\bm{\xi}}}^{M}_{t})\right\|^{2}$ to update the model parameters $\Theta$.

### v2

@@ Abstract
Modern physics simulation often involves multiple functions of interests, and traditional numerical approaches are known to be complex and computationally costly. While machine learning-based surrogate models can offer significant cost reductions, most focus on a single task, such as forward prediction, and typically lack uncertainty quantification — an essential component in many applications. To overcome these limitations, we propose Arbitrarily-Conditioned Multi-Functional Diffusion (ACM-FD), a versatile probabilistic surrogate model for multi-physics emulation. ACM-FD can perform a wide range of tasks within a single framework, including forward prediction, various inverse problems, and simulating data for entire systems or subsets of quantities conditioned on others. Specifically, we extend the standard Denoising Diffusion Probabilistic Model (DDPM) for multi-functional generation by modeling noise as Gaussian processes (GP). We propose a random-mask based, zero-regularized denoising loss to achieve flexible and robust conditional generation. We induce a Kronecker product structure in the GP covariance matrix, substantially reducing the computational cost and enabling efficient training and sampling. We demonstrate the effectiveness of ACM-FD across several fundamental multi-physics systems. The code is released at https://github.com/BayesianAIGroup/ACM-FD.
@@ 1 Introduction
Physical simulation plays a crucial role in numerous scientific and engineering applications. Traditional numerical approaches (Zienkiewicz et al., 1977; Mitchell & Griffiths, 1980), while offering strong theoretical guarantees, are often complex to implement and computationally expensive to run. In contrast, machine learning-based surrogate models — also known as emulators — are trained on simulation or measurement data and can significantly reduce computational costs, making them a promising alternative (Kennedy & O’Hagan, 2000; Razavi et al., 2012).
However, modern physical simulations often involve multiple functions of interest, such as initial and boundary conditions, solution or state functions, parameter functions, source functions, and more. Current machine learning-based surrogate models, such as neural operators (Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2023), primarily focus on a single prediction task, for instance, forward prediction of the solution function. To perform other tasks, one typically needs to retrain a surrogate model from scratch. In addition, most existing methods do not support uncertainty quantification, which is important in practice. For example, confidence intervals are important to assess the reliability of emulation results, and in inverse problems, a posterior distribution is required since such problems are often ill-posed.
@@
  • Multi-Functional Diffusion Framework: We propose a multi-functional diffusion framework based on the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). By modeling the noise as multiple Gaussian processes (GPs), we perform diffusion and denoising in functional spaces, enabling the generation of multiple functions required in multi-physics systems.
  • Innovative Denoising Loss: We introduce a denoising loss that encapsulates all possible conditional parts within the system. During training, we repeatedly sample conditional components using random masks, training the denoising network not only to restore noise for the parts to be generated, but also to predict zero values for the conditioned components. This regularization stabilizes the network and prevents excessive perturbations in the conditioned components during generation. This way, ACM-FD can flexibly generate function values conditioned on any given set of functions or quantities, allowing it to tackle a wide range of tasks, including forward prediction, inverse inference, completion, and system simulation.
  • Efficient Training and Sampling: To enable efficient training and sampling, we use a multiplicative kernel to induce a Kronecker product structure within the GP covariance matrix. By leveraging the properties of the Kronecker product and tensor algebra, we bypass the need to compute the full covariance matrix and its Cholesky decomposition, substantially reducing the training and sampling costs.
  • Experiments: We evaluated ACM-FD on four fundamental multi-physics systems, with the number of involved functions ranging from three to seven. In twenty-four prediction tasks across these systems, ACM-FD consistently achieved top-tier performance compared to state-of-the-art neural operators specifically trained for each task. Furthermore, we evaluated ACM-FD in emulating all functions jointly and function completion. The data generated by ACM-FD not only closely adheres to the governing equations but also exhibits strong diversity. Its quality is comparable to that of a diffusion model trained exclusively for unconditional generation. In parallel, ACM-FD achieves substantially higher completion accuracy than popular inpainting and interpolation methods. ACM-FD also provides superior uncertainty calibration compared to alternative approaches. Finally, a series of ablation studies confirmed the effectiveness of the individual components of our method.
@@ 2 Preliminaries
where $\widehat{\beta}_{t}=\frac{1-\widehat{\alpha}_{t-1}}{1-\widehat{\alpha}_{t}}\beta_{t}$. Accordingly, DDPM employs a neural network $\Phi$ to approximate the noises $\bm{\xi}_{t}$; see (2). The training objective is to minimize the denoising loss, $\mathbb{E}_{t}\|\Phi_{\Theta}({\bf x}_{t},t)-\bm{\xi}_{t}\|^{2}$.
@@ 3 Method
We consider a multi-physics system involving a collection of $M$ functions. These functions may represent initial and boundary conditions, source terms, parameter functions, system states and others. Our goal is to capture the underlying complex relationships between these functions and perform a wide range of prediction and simulation tasks, such as forward prediction of the system state given the initial condition, inverse inference about the systems parameters given observed states, joint simulation of all the functions, or conditional simulation of one set of functions based on another.
Many of these tasks involve mapping between functions, making neural operators (Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2023) — a recently developed class of surrogate models — an appealing approach. However, most neural operators are trained for a single prediction task. To perform a different task, one would need to train another neural operator from scratch. These models generally lack uncertainty quantification, which is important for many applications. Furthermore, these models are unable to carry out data generation.
To address these issues, we propose ACM-FD, a new generative model that can serve as a powerful and versatile probabilistic emulator for multi-physics systems.
Figure 1: The illustration of ACM-FD. The grey shaded area represents zero values introduced by the random mask. The denoising network is trained to predict zeros for the conditioned part while recovering noises for the part to be generated (i.e., unshaded noises).
@@ 3.1 Multi-Functional Diffusion
where $f_{t}(\cdot)$ is the noisy version of $f_{0}$ at step $t$, and $\xi_{t}(\cdot)$ is the noise function used to corrupt $f_{0}$. To model the noise function, one can employ an expressive stochastic process. We choose to sample each noise function from a zero-mean Gaussian process (GP) (Rasmussen & Williams, 2006),
@@
where $\kappa(\cdot,\cdot)$ is the covariance (kernel) function and ${\bf z}$ and ${\bf z}^{\prime}$ denote the input locations of the function. When $t$ becomes sufficiently large (i.e., $\widehat{\alpha}_{t}\approx 0$), we can approximately view $f_{t}\sim\mathcal{GP}(0,\kappa({\bf z},{\bf z}^{\prime}))$, meaning it effectively turns into a noise function.
The actual data is the function instance sampled at a finite set of locations, ${\bf f}_{0}=f_{0}({\mathcal{Z}})=\left(f_{0}({\bf z}_{1}),\ldots,f_{0}({\bf z}_{N})\right)^{\top}$ where ${\mathcal{Z}}=\{{\bf z}_{j}|1\leq j\leq N\}$ are the sampling locations. Therefore, we only need to apply the diffusion process (5) to ${\mathcal{Z}}$ (with the function values at all the other locations marginalized out), which yields:
@@
Comparing to (4), the only difference is that the identity matrix is replaced by the covariance matrix ${\bf K}$ at the sampling locations ${\mathcal{Z}}$. We then use a similar strategy as in DDPM. We train a neural network $\Phi_{\Theta}$ such that $\Phi_{\Theta}({\bf f}_{t},t,{\mathcal{Z}})\approx\bm{\xi}_{t}$, and substitute the neural network’s prediction for $\bm{\xi}_{t}$ in (8) to obtain $p_{\Theta}({\bf f}_{t-1}|{\bf f}_{t})$ for denoising and generation.
Next, we generalize to the multi-functional case. Given $M$ functions of interest, =$\mathcal{F}=\{f^{1}_{0}(\cdot),\ldots,f^{M}_{0}(\cdot)\}$, we employ the same diffusion process as specified in (5) and (7) to transform each function into a noise function. Next, we use a single denoising neural network to jointly recover or sample all the functions, thereby capturing the complex and strong relationships among them. Specifically, we construct a network $\Phi_{\Theta}$ that takes the step $t$, all the noisy function values sampled at $t$, and the sampling locations as the input, to predict the corresponding noises,
@@
where each ${\bf f}^{k}_{t}=f^{k}_{t}({\mathcal{Z}}_{k})$, $\bm{\xi}^{k}_{t}=\xi^{k}_{t}({\mathcal{Z}}_{k})$, $f^{k}_{t}(\cdot)$ is the noisy version of $f^{k}_{0}(\cdot)$ at step $t$, $\xi^{k}_{t}(\cdot)$ is the corresponding noise function that corrupts $f^{k}_{0}$ to yield $f^{k}_{t}$, ${\mathcal{Z}}_{k}$ are the sampling locations of $f^{k}_{0}$, and $\overline{{\mathcal{Z}}}=\{{\mathcal{Z}}_{k}\}_{k=1}^{M}$.
@@ 3.2 Arbitrarily-Conditioned Denoising Loss
Specifically, let us denote an arbitrary set of conditioned function values as ${\bf F}^{c}=\{f_{0}^{k}({\mathcal{Z}}_{k}^{c})|k\in\mathcal{C}\}$, and the target function values to be generated as ${\bf F}^{s}=\{f_{0}^{k}({\mathcal{Z}}_{k}^{s})|k\in{\mathcal{S}}\}$. Here $\mathcal{C}$ and ${\mathcal{S}}$ denote the indices of the functions to be conditioned on and those to be generated, respectively. Note that $\mathcal{C}$ and ${\mathcal{S}}$ may overlap, provided that for any $k\in\mathcal{C}\cap{\mathcal{S}}$, the sampling locations do not, i.e., ${\mathcal{Z}}_{k}^{c}\cap{\mathcal{Z}}_{k}^{s}=\emptyset$. To generate ${\bf F}^{s}$ conditioned on ${\bf F}^{c}$, we fix ${\bf F}^{c}$ in the input to the network $\Phi_{\Theta}$ as a constant, and vary only the noisy state associated with ${\bf F}^{s}$, denoted as ${\bf F}^{s}_{t}$ (at step $t$). The input to $\Phi$ consists of ${\bf F}^{c}\cup{\bf F}^{s}_{t}\cup\{t,\overline{{\mathcal{Z}}}\}$. We use the predicted noise corresponding to ${\bf F}^{s}_{t}$ to sample ${\bf F}^{s}_{t-1}$, following the denoising process. This procedure is repeated until $t=0$, resulting in a sample for ${\bf F}^{s}$. The sampling process is summarized in Algorithm 2.
@@
By varying the choices of ${\bf F}^{c}$ and ${\bf F}^{s}$, our model can perform a wide variety of data generation and prediction tasks. When ${\bf F}^{c}=\emptyset$, the model jointly samples the $M$ target functions at the specified input locations. When ${\bf F}^{c}\neq\emptyset$ and $\mathcal{C}\cap{\mathcal{S}}=\emptyset$, the model generates one set of functions conditioned on the other, supporting various forward prediction and inverse inference tasks. For instance, by setting $\mathcal{C}$ to include initial conditions or current and/or past system states and source functions, and ${\mathcal{S}}$ to the future system states, the model performs forward prediction. Alternatively, if the observed system states are included in $\mathcal{C}$ and ${\mathcal{S}}$ corresponds to initial/boundary conditions, system parameters, or past states, the model carries out inverse inference. Even for the same type of tasks, the functions and the number of functions in $\mathcal{C}$, as well as the number of sampling locations in ${\mathcal{Z}}_{k}^{c}$, can be adjusted to accommodate different levels of available information for prediction. When $\mathcal{C}\cap{\mathcal{S}}\neq\emptyset$, the model performs not only conditional sampling of other functions but also completion of the observed functions at new locations. All tasks can be executed using a single model, $\Phi_{\Theta}$, during inference, following the same sampling process. One can generate as many predictive samples as needed, naturally supporting uncertainty quantification in arbitrary context. Thereby, our method offers a versatile probabilistic surrogate for multi-physics emulation.
@@
However, to build a well-performing model, training $\Phi$ to predict all the noises as specified in (9) is far from sufficient, because it only handles unconditioned sampling. To train our model to handle all kinds of tasks, we introduce a random mask variable, =${\mathcal{H}}=\{{\bf h}^{1},\ldots,{\bf h}^{M}\}$. Each element in ${\mathcal{H}}$ is binary, indicating whether the corresponding function value is conditioned or to be sampled. If a function value is conditioned, then it should remain fixed as a constant input to $\Phi$, and correspondingly, the predicted noise for that value should be zero, since it is never corrupted by any noise. Accordingly, we introduce a new denoising loss,
@@
Here ${\bf f}^{k}_{0}$ is the original function instance sampled at the specified locations ${\mathcal{Z}}^{k}$, ${\bf f}^{k}_{t}=\sqrt{\widehat{\alpha}_{t}}{\bf f}^{k}_{0}+\sqrt{1-\widehat{\alpha}_{t}}\bm{\xi}^{k}_{t}$ is the noisy version of ${\bf f}^{k}_{0}$ at step $t$, and $\circ$ denotes the element-wise product. From (11), we can see $\widehat{{\bm{\xi}}}^{k}_{t}$ is mixed with zeros for the conditioned part and the noises for the part to generate. The zeros in $\widehat{{\bm{\xi}}}^{k}_{t}$ regularizes the network $\Phi_{\Theta}$, preventing it from producing excessively large perturbations in the conditioned part. This encourages the model to better adhere to the system’s underlying mechanism (e.g., governing equations) during generation. See Fig. 1 for an illustration.
The choice of $p({\mathcal{H}})$ is flexible. We may randomly mask individual observed values within each function and/or randomly mask entire functions. In the absence of prior knowledge or specific preferences, a natural choice is to set the masking probability to 0.5, as it yields the maximum masking variance (i.e., diversity). Our denoising loss (10) accommodates all possible conditional components for conditioned sampling as well as unconditioned sampling (i.e., ${\mathcal{H}}$ consists of all zeros), allowing our model to be trained for all kinds of tasks. We use stochastic training, which involves randomly sampling the mask ${\mathcal{H}}$, applying it to the function and noise instances as specified in (11), and computing the stochastic gradient $\nabla_{\Theta}\left\|\Phi_{\Theta}(\widehat{{{\bf f}}}^{1}_{t},\ldots,\widehat{{{\bf f}}}^{M}_{t},t,\overline{{{\mathcal{Z}}}})-(\widehat{{\bm{\xi}}}^{1}_{t},\ldots,\widehat{{\bm{\xi}}}^{M}_{t})\right\|^{2}$ to update the model parameters $\Theta$.